# Hao Shi1  Bin Xie2  Yingfei Liu2  Lin Sun4  Fengrong Liu5  Tiancai Wang2 Erjin Zhou2  Haoqiang Fan2  Xiangyu Zhang3,6  Gao Huang1🖂

### v1

@@ C.2 LIBERO
C.3 Real-world
@@
Method
Spoon
on Towel
Carrot
on Plate
Stack
Cube
Eggplant
in Basket
Avg.
Success
(a) Memory Type
Cog. Mem. 70.8 58.3 29.2 95.8 63.5
@@
(c) Memory Retrieval
w/o Timestep PE 83.3 62.5 50.0 83.3 69.8
@@
(e) Memory Update
FIFO 66.7 66.7 33.3 100.0 66.7
@@
Appendix D Additional Experimental Details
Tab. 8 provides an extended version of Tab. 6 and 6, reporting per-task success rates on SimplerEnv-Bridge for all ablation settings. Gray rows indicate the default configuration.
Appendix E Task Details
@@
E.1 Real-world Tasks
@@ E.2 SimplerEnv Tasks
Long put both OBJ1 and OBJ2 in the CONTAINER 10
@@
90 open/close CONTAINER/APPLIANCE [and put OBJ on/in it; optionally sequence with another open/close] 90
@@
E.3 LIBERO Tasks
Tab. 11 outlines the five suites of the LIBERO benchmark: Spatial, Object, Goal, Long, and LIBERO-90. LIBERO-Spatial consists of tasks where the same object must be placed across varying target positions. LIBERO-Object focuses on handling diverse objects within a fixed scene layout. LIBERO-Goal contains heterogeneous operations such as opening containers, placing objects, or turning on appliances, performed in an unchanged environment. LIBERO-Long (also called LIBERO-10) introduces ten extended tasks that require multiple sub-goals across different scenes, while LIBERO-90 expands this setting to ninety tasks, providing a substantially more challenging benchmark. In total, LIBERO offers 130 tasks in simulation with a Franka robot.
Appendix F Qualitative Results in Real-world
We present qualitative examples to complement the quantitative evaluation. Fig. 10 and 11 illustrate rollouts on long-horizon temporal tasks in real-world. Fig. 12 shows general manipulation tasks in real-world.
@@ Appendix F Qualitative Results in Real-world
Appendix G Qualitative Results in Simulation
Results on simulated environments are visualized in Fig. 13 and 14, covering both Bridge and Fractal suites. Finally, Fig. 15 provides representative trajectories on LIBERO, spanning all five suites.
@@ Appendix G Qualitative Results in Simulation
Figure 15: Qualitative results of MemoryVLA on LIBERO tasks. Representative examples include tasks from Spatial, Object, Goal, Long, and LIBERO-90 suites.

### v2

@@ D.2 LIBERO
D.3 Mikasa-Robo
For Mikasa-Robo (cherepanov2025memory), we evaluate with 100 episodes per task. Validation is performed every 1k training steps, and success rates are reported at the checkpoint with the best validation performance. As in LIBERO, we do not use any action ensemble strategy in our Mikasa-Robo experiments.
D.4 Real-world
@@
Appendix E Case Study of Memory Retrieval
To provide a direct view of how the memory mechanism functions, Fig. 10 visualizes the retrieved memory elements and their attention weights on the real-world and simulation tasks. The model consistently attends to past frames that resolve decision-relevant ambiguities absent from the current observation.
In the real-world Change Food task, after the first food item is placed aside, the current frame contains two food items on the table, making it impossible to determine from this single observation which one should be picked next. MemoryVLAtherefore attends strongly to the nearby frames reflecting the recent motion trend, as well as the last decisive frame before the ambiguity arises. In the Shell Game Touch task from Mikasa-Robo Simulation Benchmark, the robot is briefly shown the cube location before it is covered by cups. The model consistently attends to the initial revealing frames, which provide the only reliable cue for identifying the correct cup. These results demonstrate that MemoryVLAretrieves meaningful temporal cues essential for disambiguating the next action, rather than simply recalling redundant visual history.
Figure 10: Case study of memory retrieval in real-world and simulated tasks. The figure visualizes the retrieved memory elements and their attention weights on the real-world Change Food task (top) and the simulated Shell Game Touch task in Mikasa-Robo (bottom). In both settings, the model consistently attends to past frames that resolve decision-relevant ambiguities absent from the current observation.
Appendix F Data Length Statistics
Tab. 9 reports the maximum, minimum, median, and average action lengths across all task suites, including SimplerEnv Evaluation (Bridge, Fractal), LIBERO (Spatial/Object/Goal and 10/90 task suites), and both real-world general and temporal tasks. For the real-world tasks, we additionally provide filtered statistics based on a motion-magnitude threshold (translation >> 1 cm or rotation >> 0.4 rad between consecutive frames) to remove frames where the end-effector motion is small.
Table 9: Action Length Statistics across all simulation (SimplerEnv Bridge/Fractal, LIBERO Spatial/Object/Goal, LIBERO-10/90) and real-world (General, Temporal) task suites. For real-world tasks, the “Filtered” versions remove frames whose end-effector motion is negligible (translation <1<1 cm and rotation <0.4<0.4 rad).
Task Suite Max Min Median Average
SimplerEnv-Bridge / Fractal 200 80 117 119
LIBERO-Spatial / Object / Goal 270 75 131 130
LIBERO-10 / 90 505 58 144 156
Real-General (Original) 1575 281 575 575
Real-General (Filtered) 213 40 81 84
Real-Temporal (Original) 7704 412 981 1672
Real-Temporal (Filtered) 902 72 236 288
Appendix G Additional Ablation Study
For memory length, it is closely tied to the episode length of each benchmark. As shown in Tab. 9, SimplerEnv, LIBERO, and our real-world general tasks have similar episode lengths, whereas real-world temporal tasks are substantially longer, even after heavy frame filtering. This naturally motivates using a larger memory length for temporal tasks. We further conducted an ablation on a representative real-world temporal task (Clean Table & Count). A moderate memory length (256) performs best, as shown in Tab. 10.
We also added an ablation on the number of cognitive tokens. As shown in Tab. 11, increasing the count from 1 to 4 brings no performance gain. Perhaps the single 4096-dim EOS token already captures sufficient semantic information, so adding more tokens does not provide additional benefit.
In addition, we extended the ablations on fusion type (Tab. 12) and consolidation strategy (Tab. 13) to LIBERO-Long-90 and a real-world long-horizon task.
Table 10: Additional ablation on memory length for both real-world and LIBERO-90 tasks.
(a) Real-World: Clean Table & Count
Memory Length Success Rate
64 78
256 (Base) 84
512 81
(b) LIBERO-Long-90 Tasks
Memory Length Success Rate
8 94.2
16 (Base) 95.6
32 95.6
Table 11: Ablation on the Number of Cognitive Tokens. Increasing the number of cognitive tokens from 1 to 4 does not improve performance. A single 4096-dim EOS token already provides sufficient semantic capacity.
Num. Cog Token Spoon on Towel Carrot on Plate Stack Cube Eggplant in Basket Avg. Success
1 (Base) 75.0 75.0 37.5 100.0 71.9
4 79.2 66.7 37.5 95.8 69.8
Table 12: Additional ablation on memory fusion type for both real-world and LIBERO-90.
Fusion Type Clean Table & Count (Real-World) LIBERO-Long-90 Tasks
Add 78 93.8
Gate 84 95.6
Table 13: Additional ablation on memory consolidation for both real-world and LIBERO-90.
Consolidation Type Clean Table & Count (Real-World) LIBERO-Long-90 Tasks
FIFO 76 94.9
Token Merge 84 95.6
Tab. 14 provides an extended version of Tab. 7 and 7, reporting per-task success rates on SimplerEnv-Bridge for all ablation settings. Gray rows indicate the default configuration.
@@ Appendix G Additional Ablation Study
Method Spoon on Towel Carrot on Plate Stack Cube Eggplant in Basket Avg. Success
(a) Memory Type Cog. Mem. 70.8 58.3 29.2 95.8 63.5
@@
(c) Memory Retrieval w/o Timestep PE 83.3 62.5 50.0 83.3 69.8
@@
(e) Memory Update FIFO 66.7 66.7 33.3 100.0 66.7
@@
Appendix H Inference Efficiency
As shown in Tab. 15, latency, throughput, and GPU memory measurements comparing our method with the baseline on two commonly used GPUs, RTX 4090 and HGX H20. The results below are averaged over 300 runs in bfloat16.
Our method achieves a latency of 0.194 s and a throughput of 82.5 Hz on RTX 4090, corresponding to only a  3.6% increase in overhead compared to the baseline. The memory usage is 16.6 GB, which is merely +0.8 GB over the baseline. These results show that the memory module is lightweight and introduces only negligible extra cost during inference.
The overhead remains small because the memory module is deliberately lightweight: each cognitive memory entry is represented by a single cognitive token, the perceptual memory is compressed to 256 channels through a perceptual compression module, and the memory consolidation step continuously merges redundant entries. Together, these designs keep both the retrieved memory size and cross-attention cost minimal, resulting in negligible additional latency and memory usage during inference.
Table 15: Inference efficiency comparison. Latency, throughput, and GPU memory are measured over 300 runs in bfloat16 with action chunk length set to 16 on RTX 4090 and HGX H20 GPU. MemoryVLA introduces only minor overhead compared to the baseline.
Model Latency (RTX 4090) Throughput (RTX 4090) Latency (HGX H20) Throughput (HGX H20) Memory
Baseline 0.187 s 85.6 Hz 0.236 s 67.8 Hz 15.8 GB
MemoryVLA 0.194 s 82.5 Hz 0.246 s 65.0 Hz 16.6 GB
Appendix I Zero-shot Task Generalization
In addition to visual OOD tests, we have added task generalization experiments to evaluate zero shot performance on unseen task categories. As shown in Tab. 16, We use Apple To Basket as the base task and test on three unseen tasks: Eggplant To Basket, Blush To Basket, and Apple To Plate. As shown in the table below, our method achieves good zero shot task generalization.
Table 16: Zero-shot task generalization results. Apple To Basket is used as the base task, and evaluation is conducted on three unseen task categories.
Model Apple To Basket (Base Task) Eggplant To Basket (OOD Task) Blush To Basket (OOD Task) Apple To Plate (OOD Task) Avg. Success (OOD Task)
MemoryVLA 4 / 5 3 / 5 2 / 5 2 / 5 0.47
Appendix J Compare with Temporal Context VLA Baselines
As shown in Tab. 17, we have added temporal-context baselines for each benchmark, including Mikasa-Robo (cherepanov2025memory), LIBERO (liu2023libero), and SimplerEnv (li2024evaluating), to provide a fair and comprehensive comparison.
In the Mikasa-Robo benchmark, a memory-dependent manipulation benchmark, no temporal-context VLA baselines are provided in the official release. To ensure a fair comparison, we therefore reproduced CronusVLA (li2025cronusvla), a contemporaneous strong VLA model that explicitly leverages historical information by aggregating multi-frame VLM features through a sliding-window module. This provides Mikasa-Robo with a temporal-context baseline, and MemoryVLAsubstantially outperforms it across all tasks.
For the LIBERO benchmark, we include the main temporal-context VLA baselines commonly used in prior work, including TTF-VLA (liu2025ttf), TraceVLA (zheng2024tracevla), 4D-VLA (zhang20254d), CronusVLA (li2025cronusvla), and MAP-VLA (li2025map). These methods represent the strongest publicly available temporal-context VLA models that have reported results on LIBERO. MemoryVLAobtains higher average success rates than these baselines.
For the SimplerEnv benchmark, we include the temporal-context VLA baselines reported in prior work, including TraceVLA (zheng2024tracevla), RoboVLMs (liu2025towards), and CronusVLA (li2025cronusvla). These methods incorporate explicit temporal modeling, and MemoryVLAachieves higher success rates than these baselines.
Table 17: Comparison with temporal-context VLA methods across diverse benchmarks. MemoryVLA outperforms temporal-context VLA baselines across both memory-focused benchmarks and general manipulation benchmarks, including Mikasa-Robo, SimplerEnv, and LIBERO.
(a) SimplerEnv Benchmark
Method Spoon on Towel Carrot on Plate Stack Cube Eggplant in Basket Avg. Success
TraceVLA (zheng2024tracevla) 12.5 16.6 16.6 65.0 27.7
RoboVLMs (liu2025towards) 45.8 20.8 4.2 79.2 37.5
CronusVLA (li2025cronusvla) 66.7 54.2 20.8 100.0 60.4
MemoryVLA (Ours) 75.0 75.0 37.5 100.0 71.9
(b) LIBERO Benchmark
Method Spatial Object Goal Long-10 Avg. Success
TTF-VLA (liu2025ttf) 73.0 84.0 81.0 58.0 74.0
TraceVLA (zheng2024tracevla) 84.6 85.2 75.1 54.1 74.8
4D-VLA (zhang20254d) 88.9 95.2 90.9 79.1 88.6
CronusVLA (li2025cronusvla) 93.8 92.8 95.6 86.5 92.2
MAP-VLA (li2025map) 96.3 98.4 95.4 83.4 93.4
MemoryVLA (Ours) 98.4 98.4 96.4 93.4 96.7
(c) Mikasa-Robo Benchmark
Model Shell Game Touch Intercept Medium Remb. Color3 Remb. Color5 Remb. Color9 Avg. Success
CronusVLA (li2025cronusvla) 32 5 31 13 9 18.0
MemoryVLA (Ours) 88 24 44 30 20 41.2
Appendix K Task Details
@@ Appendix I Zero-shot Task Generalization
K.1 Real-world Tasks
@@
Long-10 put both OBJ1 and OBJ2 in the CONTAINER 10
@@
Long-90 open/close CONTAINER/APPLIANCE [and put OBJ on/in it; optionally sequence with another open/close] 90
@@
K.3 LIBERO Tasks
Tab. 20 outlines the five suites of the LIBERO benchmark: Spatial, Object, Goal, Long-10, and Long-90. LIBERO-Spatial consists of tasks where the same object must be placed across varying target positions. LIBERO-Object focuses on handling diverse objects within a fixed scene layout. LIBERO-Goal contains heterogeneous operations such as opening containers, placing objects, or turning on appliances, performed in an unchanged environment. Long-10 introduces ten extended tasks that require multiple sub-goals across different scenes, while Long-90 expands this setting to ninety tasks, providing a substantially more challenging benchmark. In total, LIBERO offers 130 tasks in simulation with a Franka robot.
Appendix L Visualization of Memory Dependence in Real-World Tasks
To clarify why real-world temporal tasks require memory, Fig. 11 highlights the key moments in several tasks where the correct action depends on past information rather than the current observation. Fig. 12 further provides a step-by-step example from the Change Food task, showing a case where the next action becomes ambiguous from a single frame and can only be resolved by recalling earlier steps. These examples illustrate that the stronger gains on real-robot tasks stem from their inherently memory-dependent nature.
Figure 11: Memory-dependent aspects of three real-world robotic tasks. The tasks Clean Table & Count, Guess Where, and Change Food all require tracking past events to make correct decisions, highlighting their inherent temporal dependence.
Figure 12: Step-by-step example of memory-dependent behavior. A sequence from the real-world Change Food task showing a case where the next action becomes ambiguous without recalling earlier steps.
Appendix M Qualitative Results
M.1 Real-world Evaluation
We present qualitative examples to complement the quantitative evaluation. Fig. 13 and 14 illustrate rollouts on long-horizon temporal tasks in real-world. Fig. 15 shows general manipulation tasks in real-world.
@@
M.2 Simulation Evaluation
Results on simulated environments are visualized in Fig. 16 and 17, covering both Bridge and Fractal suites. Finally, Fig. 18 provides representative trajectories on LIBERO, spanning all five suites.
@@
Figure 18: Qualitative results of MemoryVLA on LIBERO tasks. Representative examples include tasks from Spatial, Object, Goal, Long-10, and Long-90 suites.